# Breakdown of broken-symmetry approach to exchange interaction

### v1

@@ Abstract
Broken-symmetry (BS) approaches are widely employed to evaluate Heisenberg exchange parameters, primarily in combination with DFT calculations. For many magnetic materials, BS-DFT calculations give reasonable estimations of exchange parameters although systematic failures have also been reported. While the latter were attributed to deficiencies of approximate exchange-correlation functional, we prove here by treating a simple model system that the broken-symmetry methodology has serious problems. Detailed analysis clarifies the intrinsic issue with the broken-symmetry treatment of low-spin states. It shows, in particular, that the error in the BS calculation of exchange parameter scales with the degree of covalency between the magnetic and the bridging orbitals. As a possible tool to overcome this intrinsic drawback of single-determinant BS approaches, we propose their extension to a minimal multiconfigurational version.
@@ I Introduction
In the study of magnetism, understanding the exchange mechanisms and evaluating the exchange coupling parameters ($J$) are of fundamental importance [1, 2, 3, 4, 5, 6, 7]. Current quantum chemistry methodologies offer approaches for evaluating the exchange parameters in various magnetic molecules and materials of experimental interest [8, 9, 10, 11, 12]. Among them, the broken-symmetry (BS) approach [13, 14, 15] has become a standard tool, especially for calculating $J$ in large systems involving several magnetic centers. This method allows the extraction of the exchange coupling by relating the energies of high-spin and BS low-spin states to the corresponding spin configurations of the Heisenberg exchange model. This approach is universal and can be applied with various quantum chemistry methods, allowing for calculating the total energy of different electronic configurations. Most calculations of this type have been performed within density functional theory (DFT) and, to a much lesser extent, within Hartree-Fock approximation. Recently, calculations of exchange parameters with BS-coupled cluster (CC) [16], BS-$G_{0}W_{0}$ [17], and BS-self-consistent GW [18, 19] method have also been reported.
Despite its simplicity, the BS-DFT approach provides reasonable exchange parameters, often close to experimental values, for many magnetic materials [20, 9, 10, 12]. Simultaneously, many quantitative and qualitative failures have been reported and discussed [21, 22]. In particular, the strong dependence of the calculated $J$ on the exchange-correlation functionals in the DFT calculation was found [23, 24, 25, 26, 27, 22]. Although the failures have often been attributed to the limitations of quantum chemistry methods [28], they may not be the sole reason.
Here, we prove the breakdown of the broken-symmetry approach for extracting $J$ by applying it to a simple Hubbard model. We demonstrate that this breakdown originates from an artificial constraint on the multiconfigurational state imposed by the broken-symmetry determinant. The error becomes especially pronounced in the case of strong covalency between the magnetic centers and bridging ligands. To overcome this drawback, we propose a calculational scheme based on a minimal version of the multiconfigurational (MC) BS approach.
II Exact vs broken-symmetry calculations of exchange parameters
To assess the performance of the BS method, we confront its prediction for the exchange parameter $J$ with the results of exact diagonalization. To facilitate the subsequent analysis, we treated the simplest possible model that contains all the necessary physical ingredients.
@@ II Exact vs broken-symmetry calculations of exchange parameters
II.1 Generic three-site model
We consider a three-center system consisting of two half-filled magnetic orbitals (1,2) and one empty orbital ($l$) at the bridging ligand group [see Fig. 1(a)]. While looking just as H-He-H on a minimal basis, such a model can describe real magnetic materials if proper orbitals 1, 2, and $l$ are chosen [33].
@@ II.1 Generic three-site model
Here, $i$ indicates orbitals, $\sigma$ is the electron spin projection, $\hat{a}_{i\sigma}^{\dagger}$ and $\hat{a}_{i\sigma}$ are electron creation and annihilation operators in the spin-orbital $i\sigma$, respectively; $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$, $\Delta$ is the gap between metal and ligand orbital levels, and $U$ and $U^{\prime}$ are the parameters of Coulomb repulsion in the magnetic and ligand orbitals, respectively. We assume the orthonormality of atomic orbitals, $\langle i|j\rangle=\delta_{ij}$. The sign of $t$ does not influence the energy levels, hence, hereafter $t\geq 0$.
@@
Despite the simplicity, this model reproduces all essential contributions to the exchange interaction between two unpaired spins except for the spin polarization of ligands. The inclusion of direct electron transfer ($t^{\prime}$) along with the intermediate one ($t$) is indispensable for quantitative analysis of exchange interaction in many magnetic materials [33]. Besides conventional kinetic antiferromagnetic and potential ferromagnetic contributions [34, 1], it also allows us to identify the ferromagnetic kinetic exchange contribution [35, 36, 37, 38, 39]. The latter plays a vital role in complexes with strong metal-ligand covalency, such as the thiophenolate-bridged heterotrinuclear complex [LFeCoFeL]3+ [29], in which one orbital of the bridging ligand group containing the diamagnetic Co(III) is strongly hybridized with the magnetic orbitals at low-spin Fe(III) sites [36, 37, 33]. Such a generic situation [Fig. 1(b)] is met for numerous bridging groups L, the limitation to one empty or doubly occupied ligand orbital $l$ being sufficient in many cases [33]. For the bridging geometry in Fig. 1(b), $t^{\prime}$ is expected to be significantly smaller than $t$. However, for strong metal-ligand covalency, the former can be far from negligible as in iron-sulfur bridged [LFeCoFeL]3+ where $t^{\prime}=0.4|t|$ [33]. At the same time, we can easily conceive M-L-M structures where $t^{\prime}$ can be of similar magnitude with or even larger than $t$ [Fig. 1(c),(d)] 111 Precisely, $t^{\prime}$ is compared with $t^{2}/\Delta$ ($\Delta\gg|t|,|t^{\prime}|$) [33]. When they are comparable, their influence on the exchange interaction is similar. .
@@
Given the above arguments, a comparison of exact and BS calculations of $J$ based on this model is expected to be conclusive in the quest for the validity of the latter. Moreover, the identification of the domain of parameters of Eq. (1) for which the discrepancy occurs will give a direct insight into the physical reasons for its breakdown  [^2] Although the model (1) provides a potential exchange contribution when one passes from the Wannier orbitals 1, 2 and $l$ to the magnetic orbitals of Anderson’s type [88], appreciable direct potential exchange interaction can already exist for the localized orbitals 1 and 2 in many magnetic materials [33]. For simplicity, we do not include this interaction in the model (1) since it does not influence the main conclusions of this work. .
@@
Figure 3: Exact and approximate exchange parameters and the nature of the antiferromagnetic state. $J$, $J_{\text{BS}}$, and $J_{\text{MFT}}$ with respect to (a) $t^{\prime}/U$ and (b), (c), (d) $\Delta/U$. The parameters used for the simulations are written in the figure.
II.2 Exact vs BS calculations of $J$
We calculate the exchange parameter of the model (1) within the full configuration interaction (CI) method, broken-symmetry Hartree-Fock (BS-HF) approximation, and the magnetic force theorem (MFT). In the first approach, we derive the exchange parameter $J$ in the Heisenberg model,
@@ II.2 Exact vs BS calculations of $J$
by using the energy gap between the lowest high- and low-spin energy levels obtained. In this equation, $\hat{\bm{S}}_{i}$ is a spin 1/2 operator on-site $i=1,2$. See for the details of the full CI Hamiltonian matrices Appendix A.1.
To estimate the exchange parameter for Eq. (2) by the BS-HF method, we use Yamaguchi’s formula [14]:
$J_{\text{BS}}$ $=$ $\frac{2(E_{\text{F}}-E_{\text{UHF}})}{\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}-\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}},$ (3)
@@
where, $\hat{\bm{S}}=\hat{\bm{S}}_{1}+\hat{\bm{S}}_{2}$, $\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}=2$ and $\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}$ is the expectation value of $\hat{\bm{S}}^{2}$ for BS-HF wave function $|\Psi_{\text{UHF}}\rangle$. See for the detailed expressions for the BS-HF calculations Appendix A.2.2. When the BS state displays well-localized spin densities on sites, one can equally well use the Noodleman’s expression for $J_{\text{BS}}$ [13] because in this case, $\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}\approx 1$. However, the latter gradually deviates from unity with the change of $t^{\prime}$ and $\Delta$ [see Appendix A.2.2]. Accordingly, the prediction based on Noodleman’s formula deviates from the result given by Eq. (3). In the limit of large M-L covalency, when the HF instability does not occur (the BS-HF determinant coincides with a restricted HF solution) so that $\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}=0$, the Noodleman’s expression will be strongly in error while Eq. (3) still correct 333 In the ab initio calculations of complexes, contrary to the HF method, the DFT calculations partly include the electron correlation through the exchange-correlation functional, an effect more pronounced for pure than hybrid functionals [70]. This effect is accounted for by the denominator of Yamaguchi’s expression which is simply proved by the fact that in the limit of exact exchange-correlation functional, when $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS-DFT}}=0$, the corresponding Eq. (3) will correctly describe the energy difference between two states with definite spin. .
We also calculate the exchange interaction using the magnetic force theorem [43]. In this approach, we compare the energy changes of the classical spin model and the BS-HF state by the rotations of the spins. Under the rotations of the antiparallel classical spins as in Fig. 2, the total energy is
$E(\theta)=-J\bm{S}^{2}\cos 2\theta\approx-J\bm{S}^{2}\left(1-2\theta^{2}\right),$ (4)

### v2

@@ Abstract
Broken-symmetry (BS) approaches are widely employed to evaluate Heisenberg exchange parameters, primarily in combination with DFT calculations. For many magnetic materials, BS-DFT calculations give reasonable estimations of exchange parameters, although systematic failures have also been reported. While the latter were attributed to deficiencies of approximate exchange-correlation functional, we prove here by treating a simple model system that the broken-symmetry methodology has serious problems. Detailed analysis clarifies the intrinsic issue with the broken-symmetry treatment of low-spin states. It shows, in particular, that the error in the BS calculation of exchange parameter scales with the degree of covalency between the magnetic and the bridging orbitals. This is due to the constraint on the form of multiconfigurational state imposed by the BS determinant, a feature common to other single-reference methods too. As a possible tool to overcome this intrinsic drawback of single-determinant BS approaches, we propose their extension to a minimal multiconfigurational version.
@@ I Introduction
In the study of magnetism, understanding the exchange mechanisms and evaluating the exchange coupling parameters ($J$) are of fundamental importance [1, 2, 3, 4, 5, 6, 7]. Current quantum chemistry methodologies offer approaches for evaluating the exchange parameters in various magnetic molecules and materials of experimental interest [8, 9, 10, 11, 12]. Among them, the broken-symmetry (BS) approach [13, 14, 15] has become a standard tool, especially for calculating $J$ in large systems involving several magnetic centers. This method allows the extraction of the exchange coupling by relating the energies of high-spin and BS low-spin states to the corresponding spin configurations of the Heisenberg exchange model. This approach is universal and can be applied with various quantum chemistry methods allowing for calculating the total energy of different electronic configurations. Most calculations of this type have been performed within density functional theory (DFT) and, to a much lesser extent, within Hartree-Fock approximation. Recently, calculations of exchange parameters with BS-coupled cluster (CC) [16], BS-$G_{0}W_{0}$ [17], and BS-self-consistent GW [18, 19] method have also been reported.
Despite its simplicity, the BS-DFT approach provides reasonable exchange parameters, often close to experimental values, for many magnetic materials [20, 9, 10, 12]. Simultaneously, many quantitative and qualitative failures have been reported and discussed [21, 22]. In particular, the strong dependence of the calculated $J$ on the exchange-correlation functionals in the DFT calculation was found [23, 24, 25, 26, 27, 22]. Although the failures have often been attributed to the limitations of quantum chemistry methods [28], these may not be the sole reason.
Here, we prove the breakdown of the broken-symmetry approach for extracting $J$ by considering a generic three-site model. We demonstrate that this breakdown originates from an artificial constraint on the multiconfigurational state imposed by the broken-symmetry determinant. The error becomes especially pronounced in the case of strong covalency between the magnetic centers and the bridging ligands. To overcome this drawback, we propose a calculational scheme based on a minimal version of the multiconfigurational (MC) BS approach.
II Generic three-site model
To assess the performance of the BS method, we confront its prediction for the exchange parameter $J$ with the results of exact diagonalization. To facilitate the subsequent analysis, we treat the simplest possible model that contains all necessary physical ingredients.
@@ II Generic three-site model
We consider a three-center system consisting of two half-filled magnetic orbitals (1,2) and one empty orbital ($l$) at the bridging ligand group [see Fig. 1(a)]. The two particles are either electrons or holes, depending on the situation. While looking just as H-He-H on a minimal basis, such a model can describe real magnetic materials if proper orbitals 1, 2, and $l$ are chosen [33].
@@
Here, $i$ indicates orbitals, $\sigma$ is the electron spin projection, $\hat{a}_{i\sigma}^{\dagger}$ and $\hat{a}_{i\sigma}$ are electron/hole creation and annihilation operators in the spin-orbital $i\sigma$, respectively; $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$, $\Delta$ is the gap between metal and ligand orbital levels, and $U$ and $U^{\prime}$ are the parameters of Coulomb repulsion in the magnetic and ligand orbitals, respectively. We assume the orthonormality of involved orbitals, $\langle i|j\rangle=\delta_{ij}$. The sign of $t$ does not influence the energy levels; hence, hereafter, $t\geq 0$.
@@
Despite the simplicity, this model reproduces all essential contributions to the exchange interaction between two unpaired spins except for the spin polarization of ligands. The inclusion of direct electron transfer ($t^{\prime}$) along with the intermediate one ($t$) is indispensable for quantitative analysis of exchange interaction in many magnetic materials [33]. Besides conventional kinetic antiferromagnetic and potential ferromagnetic contributions [34, 1], it also allows us to identify the ferromagnetic kinetic exchange contribution [35, 36, 37, 38, 39]. The latter plays a vital role in complexes with strong metal-ligand covalency, such as the thiophenolate-bridged heterotrinuclear complex [LFeCoFeL]3+ [29], in which one orbital of the bridging ligand group containing the diamagnetic Co(III) is strongly hybridized with the magnetic orbitals at low-spin Fe(III) sites [36, 37, 33]. Such a generic situation [Fig. 1(b)] is met for numerous bridging groups L, the limitation to one empty or doubly occupied ligand orbital $l$ being sufficient in many cases [33]. For the bridging geometry in Fig. 1(b), $t^{\prime}$ is expected to be significantly smaller than $t$. However, for strong metal-ligand covalency, the former can be far from negligible as in iron-sulfur bridged [LFeCoFeL]3+ where $t^{\prime}=0.4|t|$ [33]. At the same time, we can easily conceive M-L-M structures where $t^{\prime}$ can be of similar magnitude with or even larger than $t$ [Fig. 1(c),(d)] 111 More precisely, $t^{\prime}$ should be compared with $t^{2}/\Delta$ ($\Delta\gg|t|,|t^{\prime}|$) [33]. When they are comparable, their effect on the exchange interaction is similar. .
@@
Given the above arguments, a comparison of exact and BS calculations of $J$ based on this model is expected to be conclusive in the quest for the validity of the latter. Moreover, the identification of the domain of parameters of Eq. (1) for which the discrepancy occurs will provide direct insight into the physical reasons for its breakdown  [^2] Although the model (1) provides a potential exchange contribution when one passes from the Wannier orbitals 1, 2 and $l$ to the magnetic orbitals of Anderson’s type [99], appreciable direct potential exchange interaction can already exist for the localized orbitals 1 and 2 in many magnetic materials [33]. For simplicity, we do not include this interaction in the model (1) since it does not affect the main conclusions of this work. .
@@
Figure 2: Exact and approximate exchange parameters. The exact $J$ (solid), $J_{\text{BS}}$ (dashed), $J_{\text{BS}^{\prime}}$ (dot-dashed), $J_{\text{SA}}$ (dotted), and $J_{\text{MFT}}$ (symbols) with respect to (a) $t^{\prime}/U$ and (b), (c), (d) $\Delta/U$. The parameters used for the simulations are written in the figure. (a) The blue and red indicate, respectively, $U^{\prime}/U=0$ and 1. (b)-(d) The gray, red, and blue indicate the data with different $t^{\prime}/U$. Abbreviation used: BS - Yamaguchi’s Eq. (4), BS’ - Noodleman’s Eq. (5), SA - spin symmetry adapted expression (7), MFT - magnetic force theorem (47).
III Exact vs broken-symmetry calculations of exchange parameters
We derive the Heisenberg model,
@@ III Exact vs broken-symmetry calculations of exchange parameters
from the ground ferromagnetic (high-spin) and antiferromagnetic (low-spin) energies of our three-center Hubbard model (1). In this equation, $\hat{\bm{S}}_{i}$ is a spin 1/2 operator on site $i=1,2$. We derive the exchange parameter $J$ using the full configuration interaction (CI) method (exact solution), and by using the broken-symmetry Hartree-Fock (BS-HF) approximation.
III.1 Exact diagonalization
In the first approach, we determine the exchange parameter $J$ in the Heisenberg model using the energy gap between the exact ferromagnetic (high-spin) and antiferromagnetic (low-spin) energy levels obtained from the full CI calculations:
$J=E_{\text{F}}-E_{\text{AF}}.$ (3)
See for the details of the full CI Hamiltonian matrices Appendix A.1.
Figure 2 shows the calculated exact $J$ (3) with solid lines. As previously demonstrated[33], the exchange parameter can be ferromagnetic ($J<0$) and antiferromagnetic ($J>0$) depending on the microscopic interaction parameters in Eq. (1). More information on the dependence of $J$ on the parameters of the microscopic model (1) can be found in Ref. 33.
III.2 Broken-symmetry approaches
III.2.1 Yamaguchi’s approach
To estimate the exchange parameter in Eq. (2) with the BS-HF method, we use Yamaguchi’s formula [14]:
$J_{\text{BS}}$ $=$ $\frac{2(E_{\text{F}}-E_{\text{BS}})}{\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}-\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}},$ (4)
where $\hat{\bm{S}}=\hat{\bm{S}}_{1}+\hat{\bm{S}}_{2}$, $\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}=2$, $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ is the expectation value of $\hat{\bm{S}}^{2}$ for BS-HF (unrestricted HF) wave function $|\Psi_{\text{BS}}\rangle$, and $E_{\text{F}}$ and $E_{\text{BS}}$ are the ferromagnetic (high-spin) and broken-symmetry energies, respectively. See for the detailed expressions for the BS-HF calculations Appendix A.2.2.
Figure 2(a) shows the calculated exchange parameters in function of $t^{\prime}/U$. The solid and the dashed lines are the exact $J$ obtained from the full CI calculations and the broken-symmetry $J_{\text{BS}}$, respectively. They exhibit similar behavior, whereas $J_{\text{BS}}$ are quantitatively and, under certain ranges of parameters, qualitatively different from the exact one. $J_{\text{BS}}$ tends to overestimate the ferromagnetic contribution due to the contamination of the ferromagnetic component of about 40-50 % in the BS-HF energy (see Appendix A.2.2). The spin contamination in the BS-HF wave function makes the variational parameters (molecular orbital coefficients) not fully optimal for the description of the ground antiferromagnetic state.
Moreover, the number of variational parameters for the BS-HF wave function is smaller than for the exact solution, leading to a poorer description of the antiferromagnetic state with the BS-HF wave function. For some range of $t^{\prime}/U$, one can also see that $J_{\text{BS}}$ qualitatively differs (has opposite sign) from the exact $J$. Thus for $0.03\lesssim t^{\prime}/U\lesssim 0.05$ and $0.22\lesssim t^{\prime}/U\lesssim 0.26$ for $U^{\prime}/U=1$ ($0.21\lesssim t^{\prime}/U\lesssim 0.24$ for $U^{\prime}/U=0$), the exact $J$ becomes antiferromagnetic, whereas $J_{\text{BS}}$ is ferromagnetic 333 The qualitative difference between $J$ and $J_{\text{BS}}$ appears due to the explicit treatment of the ligand atom. Within the two-site Hubbard model, the behavior of $J$ and $J_{\text{BS}}$ as function of the parameters is similar [100]. . The discrepancy is enlarged with the increase of the Coulomb repulsion on the bridging site, $U^{\prime}$, implying that the mean-field description is not adequate. This means that the static electron correlation involving explicitly ligand type configurations is crucial to derive accurate $J$.
We can see the importance of the electron correlation for the description of $J$ by modulating the metal-ligand covalency via $\Delta$. From a detailed analysis of the present model, we have demonstrated earlier that the static electron correlation effect is enhanced for small $\Delta$ [33]. Figures 2(b), (c) show that $J_{\text{BS}}$ deviates from exact $J$ when $\Delta$ is diminished and the covalency effects are enhanced. The discrepancy is further enhanced by turning on $U^{\prime}$.
III.2.2 Noodleman’s or mapping approach
When the BS state displays well-localized spin densities on sites, one can equally well use the Noodleman’s expression for the exchange parameter (further denoted $J_{\text{BS}^{\prime}}$) [13],
$J_{\text{BS}^{\prime}}$ $=2(E_{\text{F}}-E_{\text{BS}}).$ (5)
@@ III.2.2 Noodleman’s or mapping approach
It corresponds to orthogonal magnetic orbitals, implying $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}\approx 1$ in the Yamaguchi’s expression (4). However, $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ gradually deviates from unity with the change of $t^{\prime}$ and $\Delta$ [see Fig. 3]. Accordingly, the prediction based on Noodleman’s formula deviates from the result given by Eq. (4). In the limit of large M-L covalency, when the HF instability does not occur (the BS-HF determinant coincides with a restricted HF solution) so that $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}=0$, the Noodleman’s expression will be strongly in error while Eq. (4) still correct 444 In the ab initio calculations of complexes, contrary to the HF method, the DFT calculations partly include the electron correlation through the exchange-correlation functional, an effect more pronounced for pure than hybrid functionals [81]. This effect is accounted for by the denominator of Yamaguchi’s expression which is simply proved by the fact that in the limit of exact exchange-correlation functional, when $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS-DFT}}=0$, the corresponding Eq. (4) will correctly describe the energy difference between two states with definite spin. .
Figure 3: $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ with respect to $t^{\prime}/U$. The interaction parameters are the same as those for Fig. 2(a).
[…]
$\langle\uparrow,\uparrow|\hat{H}_{\text{ex}}|\uparrow,\uparrow\rangle$ $=JS_{1}S_{2},$
$\langle\uparrow,\downarrow|\hat{H}_{\text{ex}}|\uparrow,\downarrow\rangle$ $=-JS_{1}S_{2},$ (6)
so that their difference $\Delta E_{\text{map}}$ corresponds to $2JS_{1}S_{2}$. In the case of $S_{1}=S_{2}=1/2$ this gives for $J$ Eq. (5).
Figure 2 shows the calculated broken-symmetry $J_{\text{BS}^{\prime}}$ (the dot-dashed lines) within Noodleman’s formula (5). As the delocalization becomes stronger by reducing $\Delta$ [Fig. 2(b)] or by increasing $t^{\prime}$ [Fig. 2(d)], the discrepancy between the $J_{\text{BS}^{\prime}}$ and exact $J$ becomes larger. Overall, Yamaguchi’s $J_{\text{BS}}$ tends to be closer to the exact $J$ than Noodleman’s $J_{\text{BS}^{\prime}}$.
IV Alternative single-reference methods for calculating $J$
At this point it is legitimate to inquire whether other single-reference based methods allow to overcome the drawbacks of the BS approach. Below we consider three of them, the spin symmetry adapted approach (SA), the method employing the magnetic force theorem (MFT) and the spin-flip approach (SF).
IV.1 Spin symmetry adapted approach
This method replaces the BS-HF state with a low-spin wave function which preserves the spin symmetry. Projecting out the AF ($S=0$) part from the BS-HF wave function, further identified as the SA-HF wave function, $|\Psi_{\text{SA}}\rangle$ (A.2.3), we can calculate the exchange parameter in the same way as in the BS approach. Thus we evaluate the exchange parameter $J_{\text{SA}}$ by replacing $E_{\text{BS}}$ and $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ in Eq. (4) with the expectation values for $|\Psi_{\text{SA}}\rangle$:
$J_{\text{SA}}=\frac{2(E_{\text{F}}-E_{\text{SA}})}{\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}-\langle\hat{\bm{S}}^{2}\rangle_{\text{SA}}}=E_{\text{F}}-E_{\text{SA}},$ (7)
where $E_{\text{SA}}$ is the low-spin state energy. For our model (1), we have $\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}=2$ and $\langle\hat{\bm{S}}^{2}\rangle_{\text{SA}}=0$. For detailed expressions of the wave functions and energy, see Appendix A.2.3.
According to our calculations (the dotted lines in Fig. 2), the spin-adapted approach partly cures the discrepancy between the exact $J$ and $J_{\text{BS}}$ 555 In practice, similar attempts have been made using DFT, while clear improvement of the calculated $J$ for real molecules has not been seen [97]. . The agreement between $J$ and $J_{\text{SA}}$ is better for weak electron correlation on the ligand atom, i.e., for smaller $U^{\prime}/U$ [Fig. 2(a)], and larger $t^{\prime}/U$ [Fig. 2(a)] and $\Delta/U$ [Fig. 2(b)-(d)]. Although the SA-HF approach resolves the spin contamination problem in the BS approach, the SA wave function possesses smaller degrees of freedom compared with the exact wave function, which leads to the discrepancy between the exact $J$ and $J_{\text{SA}}$ when the electron correlation effects becomes stronger.
IV.2 Magnetic force theorem
Another popular method for the calculation of exchange parameters is based on the use of magnetic force theorem [47, 48, 49]. In this approach, the ground-state BS energy, minimized under constrained directions of magnetization on magnetic sites, is confronted with a classical spin model with similar directions of the corresponding spins. Under the rotations of the antiparallel classical spins from initial collinear arrangement (Fig. 4), the total energy reads
$E(\theta)=-J_{\text{MFT}}\bm{S}^{2}\cos 2\theta\approx-J_{\text{MFT}}\bm{S}^{2}\left(1-2\theta^{2}\right),$ (8)